400 people were at a baseball game. Student tickets cost $2 and adult tickets cost $3. Total ticket sales were $1050. How many tickets of each type sold?
Answer by richwmiller(17219)

s+a=400,
2*s+3*a=1050
s=400-a
2*(400-a)+3*a=1050
800-2a+3*a=1050
1*a=250
a=250
s=400-a
s=150 at $2 and a=250 at $3
check
150+250=400
2*s+3*a=1050
2*150+3*250=1050
300+750=1050
1050=1050
ok